OpenStudy (anonymous):

Simplify: 6+4=? A: 2 B: 0 C: 7 D: 10

OpenStudy (alexandervonhumboldt2):

please close this spamm

OpenStudy (anonymous):

HOW IS THIS SPAM?!?!?!?!

OpenStudy (alexandervonhumboldt2):

anyways if you ask this you are less than 13 so you need to be banned

OpenStudy (alexandervonhumboldt2):

this is not a 13+ question

OpenStudy (anonymous):

IM ON MY BROTTHERS ACCOUNT

OpenStudy (alexandervonhumboldt2):

make your own then

OpenStudy (anonymous):

HOW IS A 6+4 a BAD QUESTIONN!?!

OpenStudy (anonymous):

Just please help.

OpenStudy (muscrat123):

sersly this is too easy to be neeeding help on. for crying out loud, count on ur fingers!

OpenStudy (alexandervonhumboldt2):

6+4=1+1+1+1+1+1+1+1+1+1=10

OpenStudy (chantysquirrel1129**):

he just needs help c:

OpenStudy (chantysquirrel1129**):

a lot of help xD

OpenStudy (muscrat123):

i know but he can count on his fingers

OpenStudy (chantysquirrel1129**):

and his toes

OpenStudy (chantysquirrel1129**):

c:

OpenStudy (muscrat123):

lol

OpenStudy (chantysquirrel1129**):

and his nose c: